An engineerAn engineer is travelling for a vacation using his car. He noticed that his current speed is consuming more fuel than the average. He decided to cut his speed from 90km/h to 60km/h to "conserve fuel" thereby cutting the fuel consumption as well from 2.5 US gal/h to 1 US gal/h. How many liters of fuel will be saved in a 100-ml trip?
1
Expert's answer
2020-08-27T05:04:23-0400

1 ml = 1.61 km

100 ml = 161 km

1) At speed 90 km/h engineer's travel time will be: 161/90= 1.8 hours

Fuel consumption in this case will be: 1.8×2.5 = 4.5 gal

2) At speed 60 km/h engineer's travel time will be: 161/60=2.7 hours

Fuel consumption in this case will be: 2.7×1=2.7 gal

Therefore, wile driving at lower speed an engineer wil save: 4.5-2.7=1.8 gal of fuel.